Q- 3 g C^2H^4 gas requires oxygen gas ( O^2) for combustion which is obtained from pure Potassium Chlorate ( KClO^3). Mass of Potassium Chlorate is ____ ?
1
Expert's answer
2015-05-21T10:14:15-0400
C2H4 + 3O2 = 2CO2+2H2O
From the equation, 1 mole of C2H4 requires 3 moles of O2 for combustion.
Consequently 3 g of C2H4requires X grams of O2
X = 3·3/1 = 9 (grams of O2)

2KClO3 = 2KCl + 3O2
From the equation, 3 moles of O2 are produced from 2 moles of Potassium Chlorate,
So 9 grams of O2 areproduced from Y grams of Potassium Chlorate
Y = 9·2/3 = 6 (grams of KClO3)
Answer: 6 grams